a bullet loses its ½ of its initial velocity at 30 cm, then at what distance it comes rest

Grade:12

1 Answers

Arun
6 years ago
let initial velovity = u
from v^2 = u^2 + 2*a*s
(u/2)^2 = u^2 + 2.a.30
 – 3u^2/4 = 60.a
a = –u^2/80
now from v = u +a*t
0 = u + ( –u^2/80) * t
t = 80/u
 
now from s = u*t + ½ *a *t^2
s = 80 + ½ * ( – u^2/80) * (80/u)^2
s = 80 – 40
s = 40 cm
 
hence it will come to rest after 40 cm.